You just purchased two coins at a price of $550 each. Because one of the coins is more collectible, you believe that its value will increase at a rate of 6.9 percent per year, while you believe the second coin will only increase at 6.3 percent per year. If you are correct, how much more will the first coin be worth in 20 years?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value of :

1st coin=550*(1.069)^20

=$2088.90(Approx).

2nd coin=550*(1.063)^20

=$1866.50(Approx).

Hence excess value=2088.90-1866.50

=$222.40(Approx).